# Revisiting In-Context Learning with Long Context Language Models

### v2

@@ 4.1 LCLM-Based ICL with Noisy Examples
LCLMs can accommodate a large number of diverse ICL examples, which raises the question of the impact and risk of including noisy examples in the context. We investigate how the performance of LCLM-enabled ICL is impacted when some or all of the ICL examples are noisy. To simulate noisy examples, we modify the outputs of a subset of in-context demonstrations by replacing their outputs with outputs from other randomly selected demonstrations. As shown in Figure 4, LCLM-enabled ICL is largely robust to noise when the proportion of noisy examples is relatively low (i.e., below 25%). This observation highlights why augmented examples, even if slightly lower quality, can still enhance performance as it increases the utilization of the context window. Also, when the amount of noise exceeds this threshold, LCLMs become vulnerable to the negative effects of noise and the performance notably declines. This adverse effect is more pronounced for challenging tasks, such as the GovReport dataset in summarization and low-resource translation tasks (e.g., English to Bemba or Ewe). This is likely because LCLMs are less familiar with those tasks, and therefore rely more on learning from in-content examples.
@@ 5 Related Work
Long Context Language Models
The field of language modeling has witnessed remarkable advancements, particularly with the development of Large Language Models (LLMs) Brown et al. (2020); OpenAI (2023); Reid et al. (2024); Dubey et al. (2024). However, early LLMs were oftentimes constrained by relatively short context windows, typically handling only a few thousand tokens at a time, which limits their applicability in advanced tasks requiring broader context comprehension, such as document-level summarization or complex reasoning Koh et al. (2023); Suzgun et al. (2022). To address this, recent efforts have led to the development of Long Context Language Models (LCLMs), designed specifically to process much larger contexts, sometimes accommodating over a million tokens within a single prompt Reid et al. (2024). To mention a few, models like Longformer and BigBird Beltagy et al. (2020); Zaheer et al. (2020) incorporate sparse attention mechanisms to efficiently handle extended contexts without compromising on computational feasibility. Recent work has pushed these limits even further – for example, LongRoPE extends the the context window of LLMs to 2M tokens by interpolating their specific positional embeddings Ding et al. (2024).
@@
Early approaches in many-shot ICL have primarily focused on the paradigm shift brought by the ability to incorporate a larger number of examples within the input context Agarwal et al. (2024); Bertsch et al. (2024), without giving much consideration to example selection strategies other than the random selection. Despite their simplicity, such many-shot ICL methods have sometimes demonstrated performance comparable to fine-tuning. Also, there is a very recent work that explores retrieval strategies in many-shot ICL Bertsch et al. (2024); however, they use models with relatively limited context capacities (e.g., under 100k tokens with Llama 2), resulting in restrictions on the number of examples included and, consequently, making retrieval-based methods appear more advantageous. However, contrary to this finding, we uncover that this advantage diminishes as the context capacity increases, allowing random sampling to perform on par with more sophisticated selection methods when a large number of examples is used. Lastly, other recent efforts include establishing benchmarks for long-context ICL Lee et al. (2024a); Li et al. (2024). Unlike prior studies, our work offers a novel perspective by systematically re-evaluating traditional selection strategies in the expanded context regime and highlighting the shift from selection optimization to effectively leveraging the extensive context space in many-shot ICL, with the proposal of data augmentation for cases where the number of examples is not sufficient to populate the context capacity of LCLMs.
@@ 6 Conclusion
We explored ICL in the context of LCLMs, which the enable inclusion of significantly more examples in-context than previously possible, and investigated whether traditional sample selection strategies remain effective in these many-shot scenarios. Through extensive experiments across diverse tasks and datasets, we observed that previously favored, sophisticated sample selection techniques offer minimal to zero performance gains over simple random selection in most cases. We believe this unexpected finding suggests a potential paradigm shift in ICL research: as LCLMs allow the processing of extensive contexts, sample selection may no longer be a priority, with simpler methods proving similarly effective and more computationally efficient due to key-value caching. We also highlighted the emerging challenge of underutilized context in low-resource tasks due to limited example availability, and, to address this, proposed a data augmentation strategy, which substantially boosts ICL performance by increasing context utilization of LCLMs. Lastly, we analyzed the behavior of LCLM-enabled ICL when operating with extremely long context and in the presence of noisy examples, and found that while performance improves with added examples, it plateaus and even declines when the context becomes too long, with increased vulnerability to noise in complex tasks. This suggests promising future directions in making LCLMs more robust to lengthy context and noise examples alongside the direction of extending their context length.
@@
While this work explores the new opportunity of ICL with LCLMs, a couple of limitations can be considered. First, the computational cost associated with LCLMs remains a significant challenge, particularly for researchers and practitioners in resource-constrained settings. Second, while the proposed data augmentation method enhances context utilization of LCLMs and improves ICL performance, the quality of synthetic examples often falls short of the quality of original data. Addressing them through cost-efficient strategies for leveraging LCLMs and developing improved data augmentation techniques would be an exciting area for future work.
@@ Appendix A Prompts
We provide the prompts used for many-shot ICL on translation, summarization, and reasoning tasks in Table A and on classification tasks in Table A. Also, we provide the prompts used for synthetic data augmentation and filtering in Table A.

### v3

@@ 4.1 LCLM-Based ICL with Noisy Examples
LCLMs can accommodate a large number of diverse ICL examples, which raises the question of the impact and risk of including noisy examples in the context. We investigate how the performance of LCLM-enabled ICL is impacted when some or all of the ICL examples are noisy. To simulate noisy examples, we modify the outputs of a subset of in-context demonstrations by replacing their outputs with outputs from other randomly selected demonstrations. As shown in Figure 7, LCLM-enabled ICL is largely robust to noise when the proportion of noisy examples is relatively low (i.e., below 25%). This observation highlights why augmented examples, even if slightly lower quality, can still enhance performance as it increases the utilization of the context window. In contrast, when the amount of noise exceeds this threshold, LCLMs become vulnerable to the negative effects of noise and the performance notably declines. This adverse effect is more pronounced for challenging tasks, such as low-resource translation (e.g., English to Bemba or Ewe). This is likely because LCLMs are less familiar with those tasks, and therefore rely more on learning from in-context examples.
@@ 5 Related Work
LCLMs
The field of language modeling has witnessed remarkable advancements with Language Models (LMs) Brown et al. (2020); OpenAI (2023); Reid et al. (2024); Dubey et al. (2024). However, earlier LMs were constrained by relatively short context windows, typically handling only a few thousand tokens at a time, which limits their applicability in advanced tasks requiring broader context comprehension, such as document-level summarization or complex reasoning Koh et al. (2023); Suzgun et al. (2022). To address this, recent efforts have led to the development of LCLMs, designed to process much larger contexts, sometimes accommodating over a million tokens within a single prompt Reid et al. (2024). To mention a few, models like Longformer and BigBird Beltagy et al. (2020); Zaheer et al. (2020) incorporate sparse attention mechanisms to efficiently handle extended contexts without compromising on computational feasibility. Also, LongRoPE extends the context window of LMs to 2M tokens by interpolating their specific positional embeddings Ding et al. (2024).
@@
Early approaches in many-shot ICL have primarily focused on the paradigm shift brought by the ability to incorporate a larger number of examples in-context Agarwal et al. (2024); Bertsch et al. (2024), without giving much consideration to example selection strategies. Such many-shot ICL methods have demonstrated performance comparable to fine-tuning. Also, there is a very recent work that explores retrieval strategies in many-shot ICL Bertsch et al. (2024); however, they use models with relatively limited context capacities (e.g., under 100k tokens with Llama 2), resulting in restrictions on the number of examples included and, consequently, making retrieval-based methods appear more advantageous. However, contrary to this finding, we uncover that this advantage diminishes as the context capacity increases, allowing random sampling to perform on par with more sophisticated selection methods when a large number of examples is used. Lastly, other recent efforts include establishing benchmarks for long-context ICL Lee et al. (2024a); Li et al. (2024). Unlike prior studies, our work offers a novel perspective by systematically re-evaluating traditional selection strategies in the expanded context regime and highlighting the shift from selection optimization to effectively leveraging the extensive context space in many-shot ICL, with data augmentation.
@@ 6 Conclusion
We explored ICL in the context of LCLMs, investigating whether traditional sample selection strategies remain effective in many-shot scenarios and observing that they offer minimal to zero performance gains over simple random selection. We also highlighted the emerging challenge of underutilized context in low-resource tasks due to limited example availability, and proposed a data augmentation strategy, which substantially boosts performance by increasing context utilization of LCLMs. Lastly, we analyzed the behavior of LCLM-enabled ICL when operating with extremely long context and in the presence of noisy examples, and found that while performance improves with added examples, it plateaus and even declines when the context becomes too long, with increased vulnerability to noise in complex tasks. This suggests promising future directions in making LCLMs more robust to lengthy context and noise examples alongside the direction of extending their context length.
@@
While this work explores the new opportunity of ICL with LCLMs, a couple of limitations can be considered. First, the computational cost associated with LCLMs remains a significant challenge, particularly for researchers and practitioners in resource-constrained settings. Second, while the proposed data augmentation method enhances context utilization of LCLMs and improves ICL performance, the quality of synthetic examples often falls short of the quality of original data. Addressing them through cost-efficient strategies for leveraging LCLMs and developing improved data augmentation techniques would be an exciting area for future work. Lastly, a theoretical understanding of why LCLMs exhibit insensitivity to example selection in many-shot settings remains an open research question.
@@ Appendix A Prompts
We provide the prompts used for many-shot ICL on translation, summarization, and reasoning tasks in Table C and on classification tasks in Table C. Also, we provide the prompts used for synthetic data augmentation and filtering in Table C.
Appendix B Detailed Experimental Setup
Configuration
For all experiments, we use the default hyperparameters for Gemini and Llama.
Ratio of Augmented Data
We use original examples alongside 1,500 synthetic samples (filtered from an initial set of 3,000 examples according to their quality scores); therefore, the percentage of augmented samples varies depending on the size of the original examples in each dataset. Specifically, for the translation task where there are around 1,000 original examples, synthetic samples comprise around 60% of the total examples. For reasoning tasks (having around 100 to 150 examples), synthetic samples constitute 90-94% of the total examples. For the classification task (e.g., Banking77 dataset), with 385 original examples, synthetic samples account for around 80% of the total examples.
Appendix C Detailed Experimental Results
Results with CoT
It is worth noting that while developing the approach to better utilize many examples within the expanded context windows of LCLMs with advanced prompting techniques, such as Chain-of-Thought (CoT) Wei et al. (2022), represents an orthogonal but promising future research direction, as an initial foray into this area, we perform experiments with CoT on the reasoning task (as it may benefit from explicit step-by-step thinking procedures) and report results in Table 7. From this, we then observe that the CoT prompting strategy improves the performance on most datasets (except for Date whose performance is already high without CoT), demonstrating that there may be a potential to enhance the performance of LCLM-enabled many-shot ICL via advanced prompting.
Table 7: Results of LCLM-enabled ICL on reasoning datasets with and without Chain-of-Thought (CoT) Wei et al. (2022).
Methods Date Salient Tracking7 Web
Many-Shot ICL 0.927 0.784 0.307 0.768
Many-Shot ICL with CoT 0.918 0.810 0.334 0.771